[figure: 9.7 cm]

CORRELATED HOPPING IN INFINITE DIMENSIONS: RIGOROUS LOCAL APPROACH

A.M. Shvaika Institute for Condensed Matter Physics Natl. Acad. Sci. Ukr. 1 Svientsitskii Str., UA-79011 Lviv, Ukraine

The main idea of the Dynamical Mean Field Theory (DMFT), which is exact in the limit of infinite dimensions, is in the local (single-site) nature of the self-energy. But this statement violate for the systems with correlated hopping, when self-energy becomes unlocal [1]. In our report we present the general approach for the description of correlated hopping in infinite dimensions which is based on the expansion around the atomic limit [2,3], keeps the DMFT local ideology and allows to calculate the thermodynamical functions. As a limit cases the Falicov-Kimball model with correlated hopping and model with broken bonds (“diluted” conductor) are considered.

[1] A. Schiller, Phys. Rev. B 60, 15660 (1999)

[2] W. Metzner, Phys. Rev. B 43, 8549 (1991)

[3] A.M. Shvaika, Phys. Rev. B 62, 2358 (2000)

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : A.M. Shvaika

Address for correspondence : Institute for Condensed Matter Physics Natl. Acad. Sci. Ukr. 1 Svientsitskii Str., UA-79011 Lviv, Ukraine

Email address : ashv@icmp.lviv.ua